OpenStudy (anonymous):

2e^2x-3=8 solve:

OpenStudy (anonymous):

add 3 divide by 2 take the log divide by 2 again

OpenStudy (anonymous):

ill try

OpenStudy (anonymous):

nvm it was answered

OpenStudy (anonymous):

still dont quite understand what to do with the log

OpenStudy (anonymous):

\[e^{\xi}=A\\ \xi =\ln(A)\]

OpenStudy (anonymous):

so when you get to \[e^{2x}=\frac{11}{2}\] the next step is \[2x=\ln(\frac{11}{2})\]

OpenStudy (anonymous):

okay, thats what i did i just wasnt sure if it was right, thank you

OpenStudy (anonymous):

yw